OpenStudy (elise_a18):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. : 69 = x + one-half(x - 6) What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (elise_a18):

so Sami, you got 3 over 2 somehow and the 6 became a 3...

sammixboo (sammixboo):

OK, so instead of doing what I said before, we can making the problem like this, 69 = \(x+\dfrac{x-6}{2}\) Because 1/2(x-6) is saying find a half (divide by 2) of x-6

sammixboo (sammixboo):

This way is easier

sammixboo (sammixboo):

Understand =)

OpenStudy (elise_a18):

yep

sammixboo (sammixboo):

OK, so now we have to multiply 2 on both sides, since we are originally dividing by 2, so we have 69 x 2 = \(x+\dfrac{x-6}{2}\times2\) Now, multiplying 2 on the RIGHT side of the equation turns x to 2x, since you are multiplying it by 2, and cancels out the fraction, so we have 69 x 2 = \(2x+x-6\) Now we need to solve 69*2, which is what?

OpenStudy (elise_a18):

138

OpenStudy (elise_a18):

then subtract 6 and divide by 2? X/

OpenStudy (elise_a18):

I have no Idea X(

sammixboo (sammixboo):

Good, so we have 138 = \(2x+x-6\) Now, we need to combine the like terms 2x and x to get 3x (because you are adding in the extra x), so we have 138 = \(3x-6\) Do you know how to solve this?

OpenStudy (elise_a18):

I think so, is the answer 44?

OpenStudy (elise_a18):

I divide by 3?

sammixboo (sammixboo):

Close, but no :( You have to first add 6 on both sides, since you are originally subtracting. You have to do the oppisite 138 + 6 = \(3x-6 + 6\) 138 + 6 = \(3x\)

sammixboo (sammixboo):

So you need to add that, then divide it by 3 =)

OpenStudy (elise_a18):

Oh right, the opposite...

sammixboo (sammixboo):

Yep =)

OpenStudy (elise_a18):

48?

sammixboo (sammixboo):

Yes! Now we know the adult price. Now do you know how to find the child price from here?

OpenStudy (elise_a18):

subtract 6

sammixboo (sammixboo):

Mhm, and that is your answer

OpenStudy (elise_a18):

i believe..

OpenStudy (elise_a18):

Thank you so much Sami, my good and faithful serv......helper XD

sammixboo (sammixboo):

=.=

sammixboo (sammixboo):

<3 No problem